40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed May 15 00:40:19 2019, gautam, Update, SUS, ETMY diagnosis plan ETMY_sensorSpectra.pdf
    Reply  Wed May 15 10:57:57 2019, gautam, Update, SUS, EY chamber opened 
       Reply  Wed May 15 17:46:24 2019, gautam, Update, SUS, ETMY inspection firstContactFiber.JPGETMY_sensorSpectra.pdf
          Reply  Wed May 15 19:36:29 2019, Koji, Update, SUS, ETMY instepction 
             Reply  Thu May 16 13:07:14 2019, gautam, Update, SUS, First contact residue removal etmy_sensors.pdfetmy_BRmode.pdf
                Reply  Thu May 16 23:31:55 2019, gautam, Update, SUS, ETMY suspension characterization 6x
                   Reply  Fri May 17 17:01:08 2019, gautam, Update, SUS, ETMY suspension characterization etmy_oplevs.pdf
          Reply  Mon May 20 11:33:46 2019, gautam, Update, SUS, ITMY inspection 
             Reply  Mon May 20 17:12:57 2019, gautam, Update, SUS, ETMY LL adjustment 
                Reply  Mon May 20 22:06:07 2019, gautam, Update, SUS, ITMY also kicked 
                   Reply  Tue May 21 00:15:21 2019, gautam, Update, SUS, Main objectives of vent achieved (?) etmy_oplevs_20190520.pdfETMY_cplxTF.pdfETMY_diagComp.pdfScreen_Shot_2019-05-21_at_12.37.08_AM.png
                      Reply  Tue May 21 21:33:27 2019, gautam, Update, SUS, ETMY HR face cleaned 
                         Reply  Wed May 22 11:53:50 2019, gautam, Update, SUS, ETMY EQ stops backed out ETMY_eigenmodes.pdf
Message ID: 14629     Entry time: Tue May 21 21:33:27 2019     In reply to: 14628     Reply to this: 14630
Author: gautam 
Type: Update 
Category: SUS 
Subject: ETMY HR face cleaned 

[koji, gautam]

We executed this plan. Photos are here. Summary:

  1. Optic was EQ-stopped (face stops only)., with the OSEMs in situ. We tried to do this as evenly as possible to avoid any magnets getting stuck on OSEMs.
  2. We used the specially procured acetone from Chub to drag wipe the HR face. This was a definite improvement, we should always get the correct grade of solvents when we attempt cleaning optics.
  3. It was observed that drag-wiping did not really have the desired cleaning effect. So Koji went in with hemostat / lens tissue soaked in acetone and wiped the HR face. This improved the situation.
  4. Applied a layer of F.C. Waited for it to dry, and then peeled it off. Under the green flashlight, the optic still looks horrific - but we decided against further drag-wiping/first-contacting. If the loss is truly 50 ppm, this is totally not a show-stopper for now.
  5. Suspension cage was replaced. EQ stops were released. Bias voltages were adjusted to bring the Oplev spot back to the center of the QPD. Now a free-swinging data collection is ongoing...
The following optics were kicked:
ETMY
Tue May 21 22:58:18 PDT 2019
1242539916

So if nothing, we got to practise this new wiping technique with OSEMs in situ successfully.

Quote:
 
  1. Prepare ETMY for first contact cleaning to remove the residual piece. 
    • Drag wipe the HR surface with dehydrated acetone 
    • Apply F.C. as usual, inspect the HR face after peeling for improvement if any.
    • This will give us a chance to practise the F.C.ing with the optic EQ-stopped (moving cage etc).
ELOG V3.1.3-